OpenStudy (anonymous):

8 – 3x – x = 12 A. –1 B. –2 C. –5 D. –10

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thank you SOOO MUCH!!!

OpenStudy (anonymous):

:)

OpenStudy (shaik0124):

8-4x=12 -4x=12-8 -4x=4 x=-1 it's A i think